OpenStudy (anonymous):

in simplest form: 3 rad 20 (2 rad 5-7)

jhonyy9 (jhonyy9):

rad wann signing squareroot ?

jhonyy9 (jhonyy9):

- so than yes 3rad20 = 3rad(4*5)= 3*2rad5 =6rad5

OpenStudy (anonymous):

these are the options a: 5 rad 10 b. 8 rad 10 c. 15 rad 10 d. 75 rad 10